# Tube formula for spherically contoured random fields with subexponential marginals

### v1

@@ Abstract
Keywords: Bonferroni approximation, critical radius, Euler characteristic heuristic, regularly varying tails, simultaneous inference, sub-Gaussian random field, upper tail dependence coefficient.
@@ Introduction
The random field (1) is isotropic, that is, invariant under the orthogonal transformations $u\mapsto gu$, $g\in O(n)$ of indices $u\in\mathbb{S}^{n-1}$. We call the random field $\Xi$ in (1) a spherically contoured random field because the distribution of the random vector $\xi$ is spherically contoured. Different from the standard spherical harmonics setting on the sphere, where second-order integrability is assumed (e.g., [18]), we explicitly allow for heavy-tailed $r=\|\xi\|$ with $\mathbb{E}r^{2}=+\infty$. The random field (1) can be alternatively introduced as
@@
where $\zeta$ is the $n$-dimensional standard Gaussian vector, and $s$ is a positive random variable stochastically independent of $\zeta$. The random field $\Xi$ in (2) is a special case of elliptically contoured random fields (see [10]). In statistical applications, $s$ in (2) is an estimator of the unknown scale parameter of $\zeta$ (Section 3).
@@
Here we note that the form (3) is canonical in the following sense. Consider a random field defined by a centered Gaussian random field $(f_{t})_{t\in T}$ multiplied by an independent random variable $s^{-1}$ (e.g., $s^{-2}$ is totally skewed $\alpha/2$-stable distributed in the $\alpha$-stable sub-Gaussian random field case [2], $20<\alpha<$ 2). Assume that $f_{t}$ is unit-variance having a finite Karhunen–Loève expansion $f_{t}=\langle\phi(t),\zeta\rangle_{\mathbb{R}^{n}}$, where $\phi(t)=(\phi_{i}(t))_{1\leq i\leq n}$. Then, the supremum of such a random field over the index set $T$ is written as
@@
with $c\to+\infty$ is exponentially small as a function of $c$, and its exponential decay rate is determined by the critical radius (reach) of the index set $M$ of the random field [13, 26]. On the other hand, little is known about non-Gaussian random fields [2, 3]. We call the approximation $P_{\mathrm{tube}}$ tail-valid if $\Delta(c)\to 0$ as $c\to+\infty$, meaning that $P_{\mathrm{tube}}(c)$ and $P(c)$ are asymptotically equivalent. The tail-validity would be a desirable property for the tail probability approximation. We call the approximation $P_{\mathrm{tube}}$ not tail-valid or tail-invalid if $\Delta(c)\not\rightarrow 0$ as $c\to+\infty$.
@@
The first example concerns the Bonferroni method for multiple testing. There, the maximum test statistic $T_{\max}=\max_{1\leq i\leq N}T_{i}$ often occurs, where $T_{i}$ are test statistics in $i$th test. It is assumed that $(T_{i})_{1\leq i\leq N}$ are jointly Gaussian random variables with covariance structure $\mathrm{Var}(T_{i})=\sigma^{2}$ and $\mathrm{Corr}(T_{i},T_{j})=\rho_{ij}$, and under the null hypothesis, their means are $\mathbb{E}[T_{i}]=0$. In a typical setting of linear models, $\rho_{ij}$ is determined by the design matrix and hence known. On the other hand, $\sigma^{2}$ is an unknown variance of measurement error. When an estimator $\widehat{\sigma}$ of $\sigma$ stochastically independent of $T_{i}$’s is available, we redefine $T_{i}$ by $T_{i}/\widehat{\sigma}$ to remove the nuisance parameter $\sigma^{2}$. One possible way to obtain such an estimator $\widehat{\sigma}$ is by the replication of measurements. By choosing unit vectors $u_{i}\in\mathbb{S}^{n-1}$ so that $\langle u_{i},u_{j}\rangle=\rho_{ij}$, it holds that $(T_{i})_{1\leq i\leq N}\stackrel{{\scriptstyle d}}{{=}}(\Xi_{u_{i}})_{1\leq i\leq N}$, where $\Xi_{u}=\langle u,\zeta\rangle/s$ with $s=\widehat{\sigma}/\sigma$. Here and in what follows, “=$\stackrel{{\scriptstyle d}}{{=}}$” denotes the equality in distribution. Hence,
@@
The second example is about the simultaneous confidence band in nonlinear (curvilinear) regression [19, 17]. We consider the case where the measurement error variance $\sigma^{2}$ is unknown. Suppose that the pairs $(x_{i},y_{i})$ of explanatory variables $x_{i}\in\mathcal{X}\subset\mathbb{R}^{n}$ and response variables $y_{i}\in\mathbb{R}$, $i=1,\ldots,N$, are observed. We consider the regression model
[…]
@@
where $u\pm v$ denotes the region $(u-v,u+v)$. The threshold $c_{\gamma}$ is determined so that the event (5) holds for all $x\in\mathcal{X}$ with given probability $1-\gamma$ [16]. That is,
@@
Although the tube method is used to approximate (6) even when $\sigma^{2}$ is unknown (see, e.g., [23]), this approximation is not yet well justified.
@@
Let $\bar{\mathbb{R}}=\mathbb{R}\cup\{\pm\infty\}$ be the extended real line and ¯$\bar{\mathbb{R}}_{+}=(0,+\infty)\cup\{+\infty\}$. For two functions $f,g:\mathbb{R}_{+}\to\mathbb{R}_{+}$, we write $f(y)\sim g(y)$ as $y\to+\infty$ iff $\lim_{y\to+\infty}f(y)/g(y)=1$, and $f(y)\asymp g(y)$ as $y\to+\infty$ iff $\lim_{y\to+\infty}f(y)/g(y)\in(0,+\infty)$. Let $\sharp A$ be the cardinality of a set $A$.
@@
where $r$ and $\eta\sim\mathrm{Unif}(\mathbb{S}^{n-1})$ are independent. Therefore, the definitions of $\Xi_{u}$ in (1) and (2) are equivalent. Introduce the notation $R_{k}=\sum_{i=1}^{k}\xi_{i}^{2}$ and $\bar{R}_{n-k}=R_{n}-R_{k}$, $k=1,\ldots,n$. Notice that $R_{n}=r^{2}=\|\xi\|^{2}$.
Let $B(x;p,q)=\int\limits_{0}^{x}y^{p-1}(1-y)^{q-1}\,\mathrm{d}y$, $x\in[0,1]$ be the incomplete Beta function, where $p,\,q>0$ and $B(p,q)=B(1;p,q)$ is the usual Beta function. Denote by $B_{p,q}$ the Beta distribution with parameters $p,q>0$.
@@
An explicit formula for $\theta(x,v)$ is given in [15, Section 2.3] as
@@
Following [25], let us introduce the fibre space
@@
The relation (7) below can be proven similarly to [25, Proposition 3.2], while relation (8) can be found in [24, Theorem 2.2]:
@@
$=2\mathbb{P}\bigl{(}R_{1}\geq c^{2},R_{1}\geq\bar{R}_{n-1}\tan^{-2}\theta(u_{1},v)\bigr{)},$
@@
In the sequel, we assume that the distribution of the random variable $R_{n}=r^{2}=\|\xi\|^{2}$ may exhibit different types of tail behavior. For this purpose, we introduce a number of distribution classes, particularly heavy-tailed ones [8, 7].
@@
$F$ is called subexponential, or $F\in\mathcal{S}$, if $\overline{F*F}(x)\sim 2\bar{F}(x)$ as ${x\to+\infty}$, where $F*F$ is the convolution of $F$ with itself.
@@
$F$ is called regularly varying with index $-\gamma>-\infty$, or $F\in\mathrm{RV}_{-\gamma}$, if and only if
@@
See [27, Theorem 2] for $\mathcal{S}\subset\mathcal{L}\subset\mathcal{H}$ as well as [8] for $\mathcal{S}^{*}\subset\mathcal{S}$ and $\mathrm{RV}_{-\gamma}\subset\mathcal{S}$. ∎
The following fact is shown in [8, Lemma 3.4]:
@@
    by Karamata’s theorem (e.g., [5, Theorem 1.5.11], [7, Theorem 2.45]), which tends to infinity as $x\to+\infty$. When $\beta=1$, we have
@@
    Set $\lambda_{0}\in(0,\gamma)$. When $\beta\leq 0$, for $q(t)=\ell(t)/t^{\beta}$, it holds that $q(t)>\lambda_{0}$ for $t$ large enough, and consequently
@@
    According to [5, Theorem 1.3.1], any slowly varying function has the form
@@
    For $q(x)=\ell(x)/x^{\beta}$ with $\beta\in(0,1)$, or $(\beta,\gamma)=(1,+\infty)$, we have $xq(x)\to+\infty$, $q(x)\to 0$, $q\in\mathrm{RV}_{-\beta}$, hence by [8, Corollary 3.9 (b)], $F\in\mathcal{S}^{*}\subset\mathcal{S}$ holds. If $q(x)=\ell(x)/x$ with $\lim_{x\to+\infty}\ell(x)=\gamma<+\infty$, then $F\in\mathrm{RV}\subset\mathcal{S}$ by (iii). When $q(x)=\ell(x)/x^{\beta}$ with $\beta\leq 0$, we have $F\in\mathcal{H}^{c}\subset\mathcal{S}^{c}$.
@@
    where $K_{j}(\cdot)$ is the modified Bessel function of the second kind. It is the probability law of the product of independent $\chi_{\nu_{1}}^{2}$- and $\chi_{\nu 2}^{2}$-distributed random variables with degrees of freedom $\nu_{1},\nu_{2}>0$ (see [28]). Its tail probability function is given by
@@
$(\beta,\gamma)$ $(-1,1)$ $(0,\frac{1}{2})$ $(\frac{1}{2},\frac{1}{2})$ $(1,+\infty)$ $(1,\gamma)$
@@
Throughout this section, we deal with the spherically contoured random field $\Xi=\{\Xi_{u},u\in M\}$ defined in (1) under Assumption 1 on its index domain $M\subset\mathbb{S}^{n-1}$. Recall that $R_{n}=r^{2}=\sum_{j=1}^{n}\xi_{j}^{2}$. For its distribution function $F_{R_{n}}(x)=\mathbb{P}(R_{n}\leq x)$, $x\in\mathbb{R}_{+}$, we now assume that $F_{R_{n}}\in\mathcal{L}_{\beta,\gamma}$ with $\beta\leq 1$, $\gamma\in\bar{\mathbb{R}}_{+}$ and corresponding function $q(t)=\ell(t)t^{-\beta}$. Let the random variable $B_{k}\sim B_{\frac{k}{2},\frac{n-k}{2}}$ be Beta distributed with parameters $k/2$ and $(n-k)/2$, $k=1,\ldots,m+1$, where $m=\dim M$. Define the expression
@@
In Theorems 2 and 4, the function $r_{\beta}$ accounts for higher-order terms in the asymptotics of the relative error $\Delta(c)$ as $c\to+\infty$. Note that by [5, Theorem 1.5.2], $\ell(hv)/\ell(h)=1+o(1)$ uniformly in $v$ on any compact interval as $h\to+\infty$, which, combined with $\ell(h)=\ell_{0}(h)(1+o(1))$, yields by Karamata’s theorem (e.g., [5, Theorem 1.5.11], [7, Theorem 2.45]) that
@@
$D_{k}(\theta,c)\sim\begin{cases}\displaystyle\frac{\cos^{k-2\beta+2}\theta\sin^{n-k-2}\theta\,\exp\bigl{\{}-c^{2(1-\beta)}\ell_{0}(c^{2})g_{\beta}\left(\cos^{2}\theta\right)-r_{\beta}(c^{2},\cos^{2}\theta)\bigr{\}}}{B\bigl{(}\frac{k}{2},\frac{n-k}{2}\bigr{)}c^{2(1-\beta)}\ell_{0}(c^{2})}&\\ &\mbox{if }\beta<1\mbox{ or }(\beta,\gamma)=(1,+\infty);\,\theta>0,\\[15.00002pt] \displaystyle\frac{\Gamma\left(\frac{n-k}{2}\right)}{B\bigl{(}\frac{k}{2},\frac{n-k}{2}\bigr{)}\{c^{2(1-\beta)}\ell_{0}(c^{2})\}^{\frac{n-k}{2}}}&\mbox{if }\beta<1\mbox{ or }(\beta,\gamma)=(1,+\infty);\,\theta=0,\\[15.00002pt] \displaystyle a_{\gamma,k}\mathbb{P}\Bigl{(}\widetilde{B}_{\gamma,k}<\cos^{2}\theta\Bigr{)}&\mbox{if }\beta=1,\gamma<+\infty\end{cases}$ (20)
@@
    Suppose that ${R_{n}}\sim\chi^{2}_{n}$ with tail tail probability function (12). Setting $\beta=0$, $\ell(x)=(1/2)-(n-2)/(2x)$, $\ell_{0}(x)=1/2$, $g_{0}(y)=y^{-1}-1$, we obtain
@@
    as $c\to+\infty$, which was shown in [13, 26].
@@
    Suppose that ${R_{n}}$ is Bessel-distributed with tail probability function (13). Setting set $\beta=1/2$, $\ell(x)=(1/2)-(\nu_{1}+\nu_{2}-3)/(4\sqrt{x})$, $\ell_{0}(x)=1/2$ and $g_{1/2}(y)=2(y^{-1/2}-1)$, we obtain
@@
Let $(T_{1},\ldots,T_{N})$ be a centered Gaussian random vector with covariance matrix $\Sigma=(\rho_{ij})$ and unit variances $\rho_{ii}=1$. The density function of $(T_{1},\ldots,T_{N})$ is denoted by $p_{T}(x_{1},\ldots,x_{N};\Sigma)$. The multivariate Mills’ ratio [20] states that
@@
[…]
@@
[…]
@@
The right panel of Figure 3 displays the graph of $(\log c,\log\Delta(c))$. The vertical values $\log\Delta(c)$ are computed using the formula (29) with $c:=c\times(\sqrt{e}/3)^{\frac{1}{2}}$. The relationship between $\log c$ and $\log\Delta(c)$ seems to be almost linear with a slope of $-1$ when $c\to+\infty$. The asymptotic formula (29) for $\Delta(c)$ (in purple) approximates the exact values (in green) well. The Bonferroni approximation in this setting is tail-valid.
@@
As in Figure 3, the Bonferroni method provides a good approximation to the simulated/exact values of tail probabilities of the maximum, and the error value $\Delta(c)$ in (31) is asymptotically very close to the exact value (in green). The right panel of Figure 4 shows the graph of $(c,\log\Delta(c))$. In this scaling, it exhibits almost linear behavior as $c\to+\infty$.
Figure 5 showcases the trivariate centered Gaussian random vector $(T_{i})_{i=1,2,3}$ with $R_{3}\sim\chi_{3}^{2}$. This corresponds to the case $s\equiv 1$ in (2). Obviously, $\mathbb{E}[R_{3}]=3$. Due to tail validity of the Bonferroni approximation, its features closely resemble those displayed in Figures 3 and 4. In the right panel, the graph of $(c^{2},\log\Delta(c))$ appears almost linear asymptotically with a slope of $-2$ as $c\to+\infty$.
@@
By the property of the elliptically contoured laws [6, Lemma 2], [21, Lemmas 2.5, 2.6], $R_{k}$ and $\bar{R}_{n-k}$ have the joint distribution
@@
Case (i). Using the notation $a:=\cos^{2}\theta\in[0,1]$, $h=c^{2}$, $p=k/2$, $q=(n-k)/2$, write $D_{k}(\theta,c)$ in (14) as
@@
Let $b=c^{2(1-\beta)}\ell_{0}\bigl{(}c^{2}\bigr{)}$ as before. Evidently, assumption “$\beta<1$ or $\beta=1$, $\gamma=\lim_{h\to+\infty}\ell_{0}(h)=+\infty$” of Theorem 4 is equivalent to $b\to+\infty$ as $c\to+\infty$. When the dimension $m=0$, the right-hand side of (23) rewrites as
@@
×$\times\frac{D}{N}b^{\frac{n-3}{2}}2^{-1}\pi^{-\frac{n-1}{2}}$

### v2

@@ Abstract
Keywords: Bonferroni approximation, critical radius, Euler characteristic heuristic, multivariate Mills’ ratio, regularly varying tails, simultaneous inference, sub-Gaussian random field, upper tail dependence coefficient.
@@ Introduction
The random field (1) is isotropic, that is, invariant under the orthogonal transformations $u\mapsto gu$, $g\in O(n)$ of indices $u\in\mathbb{S}^{n-1}$. We call the random field $\Xi$ in (1) a spherically contoured random field because the distribution of the random vector $\xi$ is spherically contoured. The random field $\Xi$ in (1) is a special case of elliptically contoured random fields (see [jensen-foutz:1989]). Different from the standard spherical harmonics setting on the sphere, where second-order integrability is assumed (e.g., [marinucci-peccati:2011]), we explicitly allow for heavy-tailed $r=\|\xi\|$ with $\mathbb{E}r^{2}=+\infty$. A class of random fields defined by
@@
where $\zeta$ is an $n$-dimensional standard Gaussian vector, and $s$ is a positive random variable stochastically independent of $\zeta$, constitutes a proper subclass of the spherical contoured random field (1). Indeed, when $\|\xi\|=r$ has a compact support, $\xi$ cannot be expressed as $\zeta/s$ in (2). For other examples of $\xi$ not expressible as $\zeta/s$, see, e.g., [Gomez-etal:2008, Theorem 2.1]. In statistical applications, $s$ in (2) can be interpreted as an estimator of the unknown scale parameter of $\zeta$ (Section 3).
@@
Here we note that the form (3) is canonical in the following sense. Consider a random field defined by a centered Gaussian random field $(f_{t})_{t\in T}$ multiplied by an independent random variable $s^{-1}$ (e.g., $s^{-2}$ is totally skewed $\alpha/2$-stable distributed in the $\alpha$-stable sub-Gaussian random field case [adler-samorodnitsky-taylor:2010], $0<\alpha<2$). Assume that $f_{t}$ is unit-variance having a finite Karhunen–Loève expansion $f_{t}=\langle\phi(t),\zeta\rangle_{\mathbb{R}^{n}}$, where $\phi(t)=(\phi_{i}(t))_{1\leq i\leq n}$. Then, the supremum of such a random field over the index set $T$ is written as
@@
with $c\to+\infty$ is exponentially small as a function of $c$, and its exponential decay rate is determined by the critical radius (reach) of the index set $M$ of the random field [kuriki-takemura:2001, taylor-takemura-adler:2005]. On the other hand, little is known about non-Gaussian random fields [adler-samorodnitsky-taylor:2010, adler-samorodnitsky-taylor:2013]. We call the approximation $P_{\mathrm{tube}}$ tail-valid if $\Delta(c)\to 0$ as $c\to+\infty$, meaning that $P_{\mathrm{tube}}(c)$ and $P(c)$ are asymptotically equivalent. The tail-validity would be a desirable property for the tail probability approximation. We call the approximation $P_{\mathrm{tube}}$ not tail-valid or tail-invalid if $\Delta(c)\not\rightarrow 0$ as $c\to+\infty$.
@@
The first example concerns the Bonferroni method for multiple testing. There, the maximum test statistic $T_{\max}=\max_{1\leq i\leq N}T_{i}$ often occurs, where $T_{i}$ are test statistics in $i$th test. It is assumed that $(T_{i})_{1\leq i\leq N}$ are jointly Gaussian random variables with covariance structure $\mathrm{Var}(T_{i})=\sigma^{2}$ and $\mathrm{Corr}(T_{i},T_{j})=\rho_{ij}$, and under the null hypothesis, their means are $\mathbb{E}[T_{i}]=0$. In a typical setting of linear models, $\rho_{ij}$ is determined by the design matrix and hence known. On the other hand, $\sigma^{2}$ is an unknown variance of measurement error. When an estimator $\widehat{\sigma}$ of $\sigma$ stochastically independent of $T_{i}$’s is available, we redefine $T_{i}$ by $T_{i}/\widehat{\sigma}$ to remove the nuisance parameter $\sigma^{2}$. One possible way to obtain such an estimator $\widehat{\sigma}$ is by the replication of measurements. By choosing unit vectors $u_{i}\in\mathbb{S}^{n-1}$ so that $\langle u_{i},u_{j}\rangle=\rho_{ij}$, it holds that $(T_{i})_{1\leq i\leq N}\stackrel{{\scriptstyle d}}{{=}}(\Xi_{u_{i}})_{1\leq i\leq N}$, where $\Xi_{u}=\langle u,\zeta\rangle/s$ with $s=\widehat{\sigma}/\sigma$. Here and in what follows, $\stackrel{{\scriptstyle d}}{{=}}$” denotes the equality in distribution. Hence,
@@
The second example is about the simultaneous confidence band in nonlinear (curvilinear) regression [naiman:1986, lu-kuriki:2017]. We consider the case where the measurement error variance $\sigma^{2}$ is unknown. Suppose that the pairs $(x_{i},y_{i})$ of explanatory variables $x_{i}\in\mathcal{X}\subset\mathbb{R}^{n}$ and response variables $y_{i}\in\mathbb{R}$, $i=1,\ldots,N$, are observed. We consider the regression model
$y_{i}=b^{\top}f(x_{i})+\varepsilon_{i},\quad\varepsilon_{i}\sim\mathcal{N}\bigl(0,\sigma^{2}\bigr),\ \ i=1,\ldots,N,\quad\mbox{i.i.d.},$
@@
where $u\pm v$ denotes the region (uv,u+v)(u-v,u+v). The threshold $c_{\gamma}$ is determined so that the event (5) holds for all $x\in\mathcal{X}$ with given probability $1-\gamma$ [liu:2010]. That is,
@@
Although the tube method is used to approximate (6) even when $\sigma^{2}$ is unknown (see, e.g., [sun-loader:1994]), this approximation is not yet well justified.
@@
Let $\bar{\mathbb{R}}=\mathbb{R}\cup\{\pm\infty\}$ be the extended real line and $\bar{\mathbb{R}}_{+}=(0,+\infty)\cup\{+\infty\}$. For two functions $f,g:\mathbb{R}_{+}\to\mathbb{R}_{+}$, we write $f(y)\sim g(y)$ as $y\to+\infty$ iff $\lim_{y\to+\infty}f(y)/g(y)=1$, and $f(y)\asymp g(y)$ as $y\to+\infty$ iff $\lim_{y\to+\infty}f(y)/g(y)\in(0,+\infty)$. Let $\sharp A$ be the cardinality of a set $A$.
@@
where $r$ and $\eta\sim\mathrm{Unif}(\mathbb{S}^{n-1})$ are independent. Therefore, $\Xi_{u}$ defined in (2) is a spherically contoured random field. Introduce the notation $R_{k}=\sum_{i=1}^{k}\xi_{i}^{2}$ and $\bar{R}_{n-k}=R_{n}-R_{k}$, $k=1,\ldots,n$. Notice that $R_{n}=r^{2}=\|\xi\|^{2}$.
Let $B(x;p,q)=\int\limits_{0}^{x}y^{p-1}(1-y)^{q-1}\,\mathrm{d}y$, $x\in[0,1]$ be the incomplete Beta function, where p,q>0p,\,q>0 and $B(p,q)=B(1;p,q)$ is the usual Beta function. Denote by $B_{p,q}$ the Beta distribution with parameters $p,q>0$.
@@
An explicit formula for $\theta(x,v)$ is given in [kuriki-takemura:2009, Section 2.3] as
@@
Following [takemura-kuriki:2003], let us introduce the fibre space
@@
The relation (7) below can be proven similarly to [takemura-kuriki:2003, Proposition 3.2], while relation (8) can be found in [takemura-kuriki:2002, Theorem 2.2]:
@@
$=\frac{1}{2}\,\mathbb{P}\bigl(R_{1}\geq c^{2},R_{1}\geq\bar{R}_{n-1}\tan^{-2}\theta(u_{1},v)\bigr),$
@@
In the sequel, we assume that the distribution of the random variable $R_{n}=r^{2}=\|\xi\|^{2}$ may exhibit different types of tail behavior. For this purpose, we introduce a number of distribution classes, particularly heavy-tailed ones [goldie-kluppelberg:1998, foss-etal:2013].
@@
$F$ is called subexponential, or $F\in\mathcal{S}$, if $\overline{F*F}(x)\sim 2\bar{F}(x)$ as $x\to+\infty$}, where FFF*F is the convolution of $F$ with itself.
@@
$F$ is called regularly varying with index $\gamma>-\infty$, or $F\in\mathrm{RV}_{-\gamma}$, if and only if
@@
See [cistjakov:1964, Theorem 2] for $\mathcal{S}\subset\mathcal{L}\subset\mathcal{H}$ as well as [goldie-kluppelberg:1998] for $\mathcal{S}^{*}\subset\mathcal{S}$ and $\mathrm{RV}_{-\gamma}\subset\mathcal{S}$. ∎
The following fact is shown in [goldie-kluppelberg:1998, Lemma 3.4]:
@@
    by Karamata’s theorem (e.g., [bingham-etal:1987, Theorem 1.5.11], [foss-etal:2013, Theorem 2.45]), which tends to infinity as $x\to+\infty$. When $\beta=1$ and for any $\varepsilon>0$, we have
@@
    Set $\lambda_{0}\in(0,\gamma)$. When $\beta\leq 0$, for $q(t)=\ell(t)/t^{\beta}$, it holds that $q(t)>\lambda_{0}$ for $t$ large enough, and consequently, for $x$ large enough,
@@
    According to [bingham-etal:1987, Theorem 1.3.1], any slowly varying function has the form
@@
    For $q(x)=\ell(x)/x^{\beta}$ with $\beta\in(0,1)$, or $(\beta,\gamma)=(1,+\infty)$, we have $xq(x)\to+\infty$, $q(x)\to 0$, $q\in\mathrm{RV}_{-\beta}$, hence by [goldie-kluppelberg:1998, Corollary 3.9 (b)], $F\in\mathcal{S}^{*}\subset\mathcal{S}$ holds. If $q(x)=\ell(x)/x$ with $\lim_{x\to+\infty}\ell(x)=\gamma<+\infty$, then $F\in\mathrm{RV}\subset\mathcal{S}$ by (iii). When $q(x)=\ell(x)/x^{\beta}$ with $\beta\leq 0$, we have $F\in\mathcal{H}^{c}\subset\mathcal{S}^{c}$.
@@
    where $K_{j}(\cdot)$ is the modified Bessel function of the second kind. It is the probability law of the product of independent $\chi_{\nu_{1}}^{2}$- and $\chi_{\nu 2}^{2}$-distributed random variables with degrees of freedom $\nu_{1},\nu_{2}>0$ (see [wells-etal:1962]). Its tail probability function is given by
@@
$(\beta,\gamma)$ (1,1)(-1,1) $(0,\frac{1}{2})$ $(\frac{1}{2},\frac{1}{2})$ $(1,+\infty)$ $(1,\gamma)$
@@
Throughout this section, we deal with the spherically contoured random field $\Xi=\{\Xi_{u},u\in M\}$ defined in (1) under Assumption 1 on its index domain $M\subset\mathbb{S}^{n-1}$. Recall that $R_{n}=r^{2}=\sum_{j=1}^{n}\xi_{j}^{2}$. For its distribution function $F_{R_{n}}(x)=\mathbb{P}(R_{n}\leq x)$, $x\in\mathbb{R}_{+}$, we now assume that $F_{R_{n}}\in\mathcal{L}_{\beta,\gamma}$ with $\beta\leq 1$, $\gamma\in\bar{\mathbb{R}}_{+}$ and corresponding function $q(t)=\ell(t)t^{-\beta}$. Let the random variable $B_{k}\sim B_{\frac{k}{2},\frac{n-k}{2}}$ be Beta distributed with parameters $k/2$ and (nk)/2(n-k)/2, $k=1,\ldots,m+1$, where $m=\dim M$. Define the expression
@@
In Theorems 2 and 4, the function $r_{\beta}$ accounts for higher-order terms in the asymptotics of the relative error $\Delta(c)$ as $c\to+\infty$. Note that by [bingham-etal:1987, Theorem 1.5.2], $\ell(hv)/\ell(h)=1+o(1)$ uniformly in $v$ on any compact interval as $h\to+\infty$, which, combined with $\ell(h)=\ell_{0}(h)(1+o(1))$, yields by Karamata’s theorem (e.g., [bingham-etal:1987, Theorem 1.5.11], [foss-etal:2013, Theorem 2.45]) that
@@
$D_{k}(\theta,c)\sim\begin{cases}\displaystyle\frac{\cos^{k-2\beta+2}\theta\sin^{n-k-2}\theta\,\exp\bigl\{-c^{2(1-\beta)}\ell_{0}(c^{2})g_{\beta}\left(\cos^{2}\theta\right)-r_{\beta}(c^{2},\cos^{2}\theta)\bigr\}}{B\bigl(\frac{k}{2},\frac{n-k}{2}\bigr)c^{2(1-\beta)}\ell_{0}(c^{2})}\hskip-170.71652pt&\\ &\mbox{if }\beta<1\mbox{ or }(\beta,\gamma)=(1,+\infty);\,\theta>0,\\[15.00002pt] \displaystyle\frac{\Gamma\left(\frac{n-k}{2}\right)}{B\bigl(\frac{k}{2},\frac{n-k}{2}\bigr)\{c^{2(1-\beta)}\ell_{0}(c^{2})\}^{\frac{n-k}{2}}}&\mbox{if }\beta<1\mbox{ or }(\beta,\gamma)=(1,+\infty);\,\theta=0,\\[15.00002pt] \displaystyle a_{\gamma,k}\mathbb{P}\Bigl(\widetilde{B}_{\gamma,k}<\cos^{2}\theta\Bigr)&\mbox{if }\beta=1,\gamma<+\infty\end{cases}$ (20)
@@
    Suppose that ${R_{n}}\sim\chi^{2}_{n}$ with tail probability function (12). Setting $\beta=0$, $\ell(x)=(1/2)-(n-2)/(2x)$, $\ell_{0}(x)=1/2$, $g_{0}(y)=y^{-1}-1$, we obtain
@@
    as $c\to+\infty$, which was shown in [kuriki-takemura:2001, taylor-takemura-adler:2005].
@@
    Suppose that $R_{n}$} is Bessel-distributed with tail probability function (13). Setting $\beta=1/2$, $\ell(x)=(1/2)-(\nu_{1}+\nu_{2}-3)/(4\sqrt{x})$, $\ell_{0}(x)=1/2$ and $g_{1/2}(y)=2(y^{-1/2}-1)$, we obtain
@@
Let $(T_{1},\ldots,T_{N})$ be a centered Gaussian random vector with covariance matrix $\Sigma=(\rho_{ij})$ and unit variances $\rho_{ii}=1$. The density function of $(T_{1},\ldots,T_{N})$ is denoted by $p_{T}(x_{1},\ldots,x_{N};\Sigma)$. The multivariate Mills’ ratio [savage:1962] states that
@@
where $\mathds{1}=(1,\ldots,1)^{\top}$. For any partition $\Sigma=\begin{pmatrix}\Sigma_{11}&\Sigma_{12}\\ \Sigma_{21}&\Sigma_{22}\end{pmatrix}$, $\Sigma^{-1}-\begin{pmatrix}\Sigma_{11}^{-1}&0\\ 0&0\end{pmatrix}$ is non-negative definite, and $\mathds{1}^{\top}\Sigma^{-1}\mathds{1}\geq\mathds{1}^{\top}\Sigma_{11}^{-1}\mathds{1}$ holds. Hence, for any $k=1,\ldots,N$ we have
@@
see [hult-lindskog:2002, Definition 2.3]. The random variables $T_{i}$, i=1, 2i=1,\,2 are said to be upper tail-independent if $\lambda_{U}=0$, and upper tail-dependent, if $\lambda_{U}>0$.
@@
The right panel of Figure 3 displays the graph of $(\log c,\log\Delta(c))$. The vertical values $\log\Delta(c)$ are computed using the formula (29) with $c:=c\times(\sqrt{e}/3)^{\frac{1}{2}}$. The relationship between $\log c$ and $\log\Delta(c)$ is asymptotically linear with a slope of $(-2)\times(-\log\cos^{2}\theta_{*})=2\log(5/8)\approx-0.94$ when $c\to+\infty$. The asymptotic formula (29) for $\Delta(c)$ (in purple) approximates the exact values (in green) well. The Bonferroni approximation in this setting is tail-valid.
@@
As in Figure 3, the Bonferroni method provides a good approximation to the simulated/exact values of tail probabilities of the maximum, and the error value $\Delta(c)$ in (31) is asymptotically very close to the exact value (in green). The right panel of Figure 4 shows the graph of $(c,\log\Delta(c))$. In this scaling, it exhibits asymptotically linear behavior with a slope of $(\sec\theta_{*}-1)\times\sqrt{4}=-2\sqrt{8/5}+2\approx-0.53$ as $c\to+\infty$.
Figure 5 showcases the trivariate centered Gaussian random vector $(T_{i})_{i=1,2,3}$ with $R_{3}\sim\chi_{3}^{2}$. This corresponds to the case $s\equiv 1$ in (2). Obviously, $\mathbb{E}[R_{3}]=3$. Due to tail validity of the Bonferroni approximation, its features closely resemble those displayed in Figures 3 and 4. In the right panel, the graph of $(c^{2},\log\Delta(c))$ appears almost linear asymptotically with a slope of $(1/2)\tan^{2}\theta_{*}=-3/10$ as $c\to+\infty$.
@@
By the property of the elliptically contoured laws [cambanis-etal:1981, Lemma 2], [schmidt:2002, Lemmas 2.5, 2.6], $R_{k}$ and $\bar{R}_{n-k}$ have the joint distribution
@@
Case (i). Using the notation $a:=\cos^{2}\theta\in[0,1]$, $h=c^{2}$, $p=k/2$, q=(nk)/2q=(n-k)/2, write $D_{k}(\theta,c)$ in (14) as
@@
Let $b=c^{2(1-\beta)}\ell_{0}\bigl(c^{2}\bigr)$ as before. Evidently, assumption $\beta<1$ or $\beta=1$, $\gamma=\lim_{h\to+\infty}\ell_{0}(h)=+\infty$” of Theorem 4 is equivalent to $b\to+\infty$ as $c\to+\infty$. When the dimension $m=0$, the right-hand side of (23) rewrites as
@@
$\times\frac{D}{N}b^{\frac{n-3}{2}}2^{-1}\pi^{-\frac{n-1}{2}}$